OpenStudy (anonymous):

Set up, but do not evaluate, the integral which gives the volume when the region bounded by the curves y = Ln(x), y = 2, and x = 1 is revolved around the line y = −2.

OpenStudy (irishboy123):

drawing it is always a good idea, especially if you are a visual thinker; but it is not obligatory.....

OpenStudy (anonymous):

@IrishBoy123 I have graphed it but I am not sure which part I am trying to find the area of. I feel like he boundaries are mixed.

OpenStudy (anonymous):

Please check for me.

OpenStudy (anonymous):

@bibby are you aware of the topic?

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

Saw you browsing the calc section so I thought you knew.

OpenStudy (anonymous):

@timo86m would you happen to know?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Can you draw it?

OpenStudy (anonymous):

Do you not see the attachment?

OpenStudy (anonymous):

Okay... no need to get so edgy :(

OpenStudy (anonymous):

Didn't mean it like that, I thought you weren't able to see it, I was gonna reupload for you :)

OpenStudy (anonymous):

No wonder you aren't asked to evaluate... Well, anyway, you got the curves right, you just shaded the wrong part ^^

OpenStudy (anonymous):

That's what I wanted to know, which part I am looking for

OpenStudy (anonymous):

So I am basically looking for the part on top of it correct?

OpenStudy (anonymous):

If so then it would be \[2-\ln(x)\]

OpenStudy (anonymous):

Yeah... what you shaded was bounded by y=-2 and not y=2 See how the shaded part doesn't part the y=2 line? ^^ It should be something that looks like this:

OpenStudy (anonymous):

Yes, exactly ^^

OpenStudy (anonymous):

Just shaded that.

OpenStudy (anonymous):

So, no more problems, then? No problem setting up the bloody integral? :D

OpenStudy (anonymous):

So it should be \[\int\limits_{0}^{2} \pi (2-\ln(x))^2\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Since y=2 is at the top

OpenStudy (anonymous):

No ^^ The limits on the integral, they should go from 0 to 1 only, as does indeed run from the y-axis (x=0) to the line x=1

OpenStudy (anonymous):

Oops, misread that. I thought x was equal to 2

OpenStudy (anonymous):

:) And don't forget the dx at the end of the integral.

OpenStudy (anonymous):

So integral setup right (aside from no dx)?

OpenStudy (anonymous):

There's just one thing that bothers me... it may well be the reason you're not asked to evaluate the integral, but... y=ln(x) is not DEFINED at x=0 as in... there is no ln(0) So can we really have a limit from 0 to 1? D:

OpenStudy (anonymous):

Yeah, I noticed that too but since I'm not solving it, I ignored it.

OpenStudy (anonymous):

@PeterPan can you look over some of my other problems?

OpenStudy (anonymous):

Sure ^^

OpenStudy (anonymous):

Tagged you. Thanks